# Multi-Task Models Adversarial Attacks

### v3

@@ Introduction
Optimization Solution via Relaxed LP: GB-MTA identifies by first addressing a relaxed Linear Programming (LP) problem and then rounding the resulting solution to obtain an integer solution. In the first step, the LP relaxation will remove the requirement of integer values, i.e. , allowing them to be any real value instead. To solve the relaxed LP, we calculate the derivative of the objective function with respect to the variable as follows,
@@
In other words, starts from the original state, that is, a zero matrix, and is updated along the direction of to maximize in the LP relaxation. Then in the second step, GB-MTA reintroduces the integer constraint, and the solution for the original ILP in Eq. 14 can be obtained by performing a rounding operation: .
@@
The optimal suggests that the effective attack direction for multi-task models should be the sum of each task’s gradients dynamically weighted by its loss value. GB-MTA mitigates the dominating task issue in Total by dynamically balancing the gradients across tasks and avoids the limited transferability problem in Single-x by optimizing over all tasks simultaneously.
@@
Integrating GB-MTA with Existing Attacks: Integrating GB-MTA with any existing attack can easily be accomplished, even for more advanced methods such as APGD (Croce and Hein 2020b). The key operation is to substitute the single task gradient, i.e., , with the balanced multi-task counterpart, i.e., . To illustrate. we provide the pseudocode for APGD integrated with GB-MTA in Algorithm C. The black text corresponds to the original APGD algorithm (Croce and Hein 2020b) and the part changed for GB-MTA is colored blue. Notice that on top of the key design shown in lines 2 and 11, we also change the absolute loss value used in the original APGD to the relative loss value sum over all the tasks in lines 4 and 14 to align with the MTL scenario.
Algorithm 1 GB-MTA-APGD
Experiments
Experimental Settings
@@
Datasets and Tasks: We use two popular datasets in multi-task learning (MTL), NYUv2 (Silberman et al. 2012) and Tiny-Taskonomy (Zamir et al. 2018). The NYUv2 dataset consists of RGB-D indoor scenes and three tasks, 40-class semantic segmentation, depth estimation, and surface normal prediction. Tiny-Taskonomy contains RGB indoor images, and its five representative tasks are semantic segmentation, surface normal prediction, depth estimation, keypoint detection, and edge detection.
@@
Evaluation Metrics and Loss Functions: Semantic segmentation uses a pixel-wise cross-entropy loss for each predicted class label. Surface normal prediction uses the inverse of cosine similarity between the normalized prediction and ground truth. All other tasks use the loss. Many tasks have distinct evaluation metrics with various scales. Hence, it is crucial to assess task performance in an equitable manner. Compounding the problem of different scales is the fact that some metrics are higher-the-better (e.g., accuracy, mean of intersection over union), while others are lower-the-better (e.g., distance, error). To address these issues and fairly measure the success of various attacks, we formulate a multi-task attack metric, Average Relative Performance (ARP):
@@
Multi-Task Models: We evaluate branched multi-task models from TreeMTL (Zhang, Liu, and Guan 2022b) using two backbone architectures: Deeplab-ResNet34 (Chen et al. 2017a) and MobileNetV2 (Sandler et al. 2018). We randomly sampled and trained 25 models with Deeplab-ResNet34 and 20 with MobileNetV2 for the NYUv2 dataset. For the Tiny-Taskonomy dataset, we sampled 15 models with Deeplab-ResNet34. These models cover a range of task sharing configurations from the all-shared models to those comprising an ensemble of independent single-task models.
Counterparts for Comparison: We compare GB-MTA to two types of baselines. The first type is the naïve multi-task attacks that repurpose existing single-task white-box attacks to multi-task models. It includes Total, SignTotal, and Single-x. We compare these baselines with GB-MTA by integrating them with three different single-task white-box attack methods, FGSM (Goodfellow, Shlens, and Szegedy 2014), PGD (Madry et al. 2018), and APGD (Croce and Hein 2020b). The second type of baseline is a multi-model attack method called Auto-SAGE, which is designed to attack multiple independent DNNs but can be directly applied to attack multi-task models. To compare fairly with Auto-SAGE, we integrate GB-MTA with Auto-SAGE instead of other single-task attacks.
Results on Attack Performance
Model Index | #Params (M) | PGD | AutoAttack | Auto-SAGE | |||||||||||
Segm | Norm | Dept | Total | SignTotal | GB-MTA | Segm | Norm | Dept | Total | SignTotal | GB-MTA | Baseline | GB-MTA | ||
@@
Model Index | #Params (M) | PGD | AutoAttack | Auto-SAGE | |||||||||||||
Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Segm | Norm | Dept | Keyp | Edge | Total | GB-MTA | Baseline | GB-MTA | ||
@@
This subsection compares GB-MTA with baselines on their effectiveness in attacking multi-task models. Tables 1 and 2 compare the attack performance of the baselines and GB-MTA at the model level on NYUv2 and Tiny-Taskonomy respectively at . Overall, GB-MTA integrated with PGD and APGD have the highest ARP in 7 out of the 8 models on NYUv2, and in 6 out of the 8 models on Tiny-Taskonomy.
@@
GB-MTA outperforms baselines Total, SignTotal, and Single-x in attack performance because it alleviates the baselines’ limitations discussed in Attack Framework. Single-x achieves limited attack effectiveness on tasks that are not the attack target x; Total shares a similar pattern of attack effectiveness across tasks as Single-Segm due to the issue of gradient dominance, making it less effective in attacking all tasks simultaneously. In contrast, GB-MTA dynamically balances the attack directions for all tasks, making it more threatening for systems that require high robustness. This is also the reason why GB-MTA outperforms the multi-model attack approach AutoSAGE — GB-MTA balances the gradients in AutoSAGE.
@@
We further vary the maximum perturbation bound from 1 to 16 and report the average of the ARP over all multi-task models with diverse architectures in Figure 2 (Deeplab-ResNet34) and Figure 5 (MobileNet). We denote the average ARP as the overall ARP. GB-MTA is the best-performing attack in almost all and dataset settings, and when integrated with any of the three existing single-task (i.e., FGSM, PGD, APGD) or the multi-model attack (i.e., Auto-SAGE) approaches. There are a few cases, where for , GB-MTA and Total converge or perform almost identically. This is because at higher values, the magnitude of the noise becomes larger (and thus more visible) and all attacks become more effective.
Results on Attack Transferability
RQ3.1: Does task sharing increase the robustness of multi-task models to adversarial attacks? Existing literature in multi-task learning determines the appropriate level of parameter sharing with a focus on optimizing task accuracy. The results reported in this section, however, reveal a fundamental trade-off between the improvement in task accuracy due to positive task interactions and the increased vulnerability to adversarial attack due to the greater transferability of attacks from parameter sharing. We observe that a higher degree of parameter sharing between correlated tasks is associated with increased attack transferability.
We first define attack transferability in the MTL context. When using an attack method like Single-x, which attacks only one task in a multi-task model, the targeted task x, would be impacted the most. Referring to Figure 1, we can see that this occurs, since the highest bar for each of the Single-x attacks is the targeted task x. Therefore, we consider the degradation of the performance of the targeted task x, represented by ARP-x, as the upper bound for the attack effectiveness of Single-x. Since adversarial examples designed to attack one task may also be adversarial (e.g. misclassified) for another task, we can quantify this phenomenon in the MTL domain. Let x represent a task under attack, and y represent another task. The transferability of attack Single-x is
where represents the number of tasks. In short, we measure transferability as a performance degradation ratio. This ratio is the performance of the attack on all non-targeted tasks versus the performance of the attack on the targeted task.
Figure 4 illustrates the relationship between the levels of task sharing and attack transferability, showcasing the results for six multi-task models attacked by APGD Single-x variants. These models represent six levels of parameter sharing, ranging from all-share (AS/5L), where all layers of the backbone model (5L) are shared, to independent models (IND/0L), where no layers are shared. We observe a positive correlation between the degree of task sharing and the transferability of attacks in the three Single-x variants. As the level of task sharing increases, attack transferability also increases, from to () for Single-Dept, to () for Single-Segm, and to () for Single-Norm. These findings suggest a trade-off in multi-task model design: While sharing more parameters among related tasks can enhance task accuracy, it also amplifies attack transferability, thereby reducing model robustness, even when facing single task attacks. This underlines the importance of balancing accuracy and robustness in multi-task model design.
Attack on Adversarially Trained Multi-Task Models
@@
RQ3.2: Does adversarial training increase MTL model robustness to adversarial attacks? A common single task strategy to defend against adversarial attacks is to leverage adversarial training (Madry et al. 2018; Bai et al. 2021; Zhang et al. 2020). This defense involves generating adversarial samples and using them as part of the dataset during training. This typically results in reduced model performance on clean inputs, but increased robustness to adversarial attacks. To the best of our knowledge, no existing work has dealt with the open question of whether adversarial training can be applied effectively to multi-task models.
Adv. Train | Single-Segm | Single-Norm | Single-Dept | Total | GB-MTA |
@@
We adopt the single task Friendly Adversarial Training (FAT) (Zhang et al. 2020) to MTL. FAT is a PGD-based adversarial training method. In order to apply this defense effectively to MTL, we modify the default PGD to GB-MTA-PGD to use adversarial examples generated from multiple tasks as opposed to just single task adversarial examples. Our new defense is coined FAT GB-MTA. We also train multi-task models using FAT with the naïve MTL attacks Single and Total. After adversarial training, we re-evaluate the multi-task attacks to assess both model robustness and attack effectiveness.
@@
Table 3 reports the ARP of 30 cases, i.e., (without adversarial training + 5 adversarially trained models) 5 attack methods. We make two main observations. First, the notable reduction in ARP of attack methods (e.g., from to when attacking with Single-Dept) demonstrates that models enhanced with adversarial training are substantially more robust than the model without such training. Second, from the perspective of attack performance, GB-MTA is still the most effective attack method, consistently outperforming the Single-x and Total baselines by up to .
Conclusion
@@
In this work, we make significant developments in multi-task learning (MTL) security through novel attack design, empirical exploration with multi-task models and robustness analyses. We first analyzed naïve adaptions of single task white-box attacks to the MTL domain and experimentally demonstrated their ineffectiveness. We then developed a new framework, Dynamic Gradient Balancing Multi-task Attack (GB-MTA), to effectively attack all tasks in multi-task models. On models trained on the NYUv2 and Tiny-Taskonomy datasets, GB-MTA achieves the highest overall attack strength and is the strongest attack on 6 out of 8 models for both datasets. We further analyzed the adversarial transferability of MTL adversarial examples and discovered a new phenomenon: Task sharing can lead to increased adversarial transferability. Lastly, from the defense side, we adversarially trained multi-task models in a new approach, which we coined FAT GB-MTA. GB-MTA is the most effective attack, even on these multi-task models, showcasing its effectiveness as a benchmark for analyzing future MTL defenses.
Acknowledgement
@@
Appendix A Related Works
@@
Adversarial Attacks. Adversarial attacks fall into two main categories: white-box and black-box attacks (Mahmood et al. 2021b). In white-box attacks, an attacker has access to the target model’s internal information, enabling direct gradient extraction and adversarial example generation (Carlini et al. 2019). On the contrary, in black-box attacks, an attacker has limited model knowledge and uses alternative information sources (Chen et al. 2017b; Zhou et al. 2020) to create adversarial examples. This paper focuses on white-box attacks, as they are generally more effective compared to black-box attacks (Croce and Hein 2020b; Wang et al. 2022).
@@
In recent years, many white-box attack techniques have been developed. Fast Gradient Sign Method (FGSM) (Goodfellow, Shlens, and Szegedy 2014) generates adversarial examples by introducing non-random noise in the gradient direction of the loss function. Projected Gradient Descent (PGD) (Madry et al. 2018) and Momentum Iterative Method (MIM) (Dong et al. 2018) improve FGSM by generating adversarial samples in an iterative process. Later, Croce and Hein (Croce and Hein 2020b) proposed Auto Projected Gradient Descent (APGD) with adaptive step size and combined it with two complementary attacks (Croce and Hein 2020a; Andriushchenko et al. 2020), developing the ensemble method APGD that outperforms existing methods on diverse benchmark datasets. In addition to gradient-based strategies, alternative methods have emerged. For instance, the Backward Pass Differentiable Approximation (BPDA) (Athalye, Carlini, and Wagner 2018) accommodates non-differentiable functions, while the Carlini and Wagner () attack (Carlini and Wagner 2017) perturbs images with minimal delta to misclassify them.
@@
Multi-Task Learning. In Multi-Task Learning (MTL), researchers develop memory and computation-efficient multi-task models that simultaneously address multiple tasks (Ruder 2017; Yao et al. 2020). The main challenge lies in determining the parameters to share across tasks to optimize both resource efficiency and task accuracy. This has led to an abundance of multi-task model architectures designed either manually (Huang et al. 2015; Jou and Chang 2016; Dvornik et al. 2017; Kokkinos 2017; Ranjan, Patel, and Chellappa 2017) or automatically (Sun et al. 2019; Zhang, Liu, and Guan 2022b; Ahn, Kim, and Oh 2019; Guo, Lee, and Ulbricht 2020; Zhang, Liu, and Guan 2022a). This paper focuses on attacking branched multi-task models, which are the most representative in the MTL literature. Besides, branched multi-task models have a wide range of sharing patterns across tasks, facilitating the study of the relationship between the robustness of multi-task models and their sharing patterns.
@@
Robustness of Multi-Task Models. Few studies have examined the robustness of the model in MTL settings. A pioneering study (Mao et al. 2020) pointed out that the adversarial robustness of deep neural networks increases as the number of tasks increases. Subsequent research (Klingner, Bar, and Fingscheidt 2020; Ghamizi et al. 2022) further emphasizes the importance of selecting suitable tasks for joint learning to create more robust models. While these studies offer intriguing insights, they do not specifically propose adversarial attack methods for multi-task models. Regarding attacks on multi-task models, MTA (Guo et al. 2020) tries to develop attacks in the MTL setting, however, the generated adversarial samples are task-specific and thus fail to attack all the tasks simultaneously. Some other work (Gurulingan, Arani, and Zonooz 2021; Sobh et al. 2021) attempted to attack the multi-task model by generating adversarial examples for each image while attacking one task at a time.
Appendix B MTL Attack Optimization Approximation

### v4

@@ 3.2 Optimization Solution via Relaxed LP
DGBA identifies by first addressing a relaxed Linear Programming (LP) problem based on Eq. 8 and then rounding the resulting solution to obtain an integer solution. In the first step, the LP relaxation will remove the requirement of integer values, i.e. , allowing them to be any real value instead. To solve the relaxed LP, we calculate the derivative of the objective function with respect to the variable as follows,
@@
In other words, starts from the original state, that is, a zero matrix, and is updated along the direction of to maximize in the LP relaxation. Then in the second step, DGBA reintroduces the integer constraint, and the solution for the original ILP in Eq. 8 can be obtained by performing a rounding operation: .
@@
The optimal suggests that the effective attack direction for multi-task models should be the sum of each task’s gradients dynamically weighted by its loss value. DGBA mitigates the dominating task issue in Total by dynamically balancing the gradients across tasks and avoids the limited transferability problem in Single-x by optimizing over all tasks simultaneously.
3.3 Integrating DGBA with Existing Attacks
DGBA can be easily integrated with any existing white-box single-task attack. The key operation is to substitute the single task gradient, i.e., , with the balanced multi-task counterpart, i.e., .
@@ 3.3 Integrating DGBA with Existing Attacks
To illustrate. we provide the pseudocode for APGD integrated with DGBA in Algorithm 3.3. We color the adaptation made by DGBA in blue. Notice that on top of the key design shown in lines 2 and 11, we also change the absolute loss value used in the original APGD to the relative loss value sum over all the tasks in lines 4 and 14 to align with the MTL scenario. This integration procedure generalizes to any gradient-based white-box single-task attack, including multi-model attacks such as Auto-SAGA. We describe the integration of DGBA with Auto-SAGA in D.
Algorithm 1 DGBA-APGD
4 Experiments
RQ2: Can adversarial attacks be designed to simultaneously attack all tasks in a multi-task model (empirically)? This section answers RQ2 by conducting experiments on different multi-task models with being attacked by DGBA to validate its effectiveness on attacking all tasks at once.
4.1 Experimental Settings
@@
Datasets and Tasks: We use two popular datasets in MTL, NYUv2 [silberman2012indoor] and Tiny-Taskonomy [zamir2018taskonomy]. The NYUv2 dataset consists of RGB-D indoor scenes and three tasks, semantic segmentation, depth estimation, and surface normal prediction. Tiny-Taskonomy contains RGB indoor images, and its five representative tasks are semantic segmentation, surface normal prediction, depth estimation, keypoint detection, and edge detection.
@@
Evaluation Metrics and Loss Functions: Semantic segmentation uses a pixel-wise cross-entropy loss for each predicted class label. Surface normal prediction uses the inverse of cosine similarity between the normalized prediction and ground truth. All other tasks use the loss. Detailed metrics for each specific task are given in E.1. Many tasks have distinct evaluation metrics with various scales. Compounding the problem of different scales is the fact that some metrics are higher-the-better (e.g., accuracy, mean of intersection over union), while others are lower-the-better (e.g., distance, error). To address the issues and fairly measure the success of attacks, we formulate a multi-task attack metric, Average Relative Performance (ARP):
@@
Multi-Task Models: We evaluate branched multi-task models from TreeMTL [zhang2022tree] using two backbone architectures: Deeplab-ResNet34 [chen2017deeplab] and MobileNetV2 [sandler2018mobilenetv2]. We randomly sampled and trained 25 models with Deeplab-ResNet34 and 20 with MobileNetV2 for the NYUv2 dataset. For the Tiny-Taskonomy dataset, we sampled 15 models with Deeplab-ResNet34. These models cover a range of parameter sharing configurations from the all-shared models to those comprising an ensemble of independent single-task models, represented by various model indexes. We provide the detailed model architectures and their corresponding model indexes in E.7.
Counterparts for Comparison: We compare DGBA to three types of baselines. (i) The first type is the adapted multi-task attacks that repurpose existing single-task white-box attacks to multi-task models. It includes Total, SignTotal, and Single-x. We compare these baselines with DGBA by integrating them with three different single-task white-box attack methods, FGSM [goodfellow2014explaining], PGD [madry2018towards], and APGD [croce2020reliable]. (ii) The second type of baseline is a multi-model attack method called Auto-SAGA [xu2022securing], which is designed to attack multiple independent DNNs but can be directly applied to attack multi-task models. We integrate DGBA with Auto-SAGA to fairly compare with the original Auto-SAGA. (iii) The third type is a multi-task model attack approach called WGD [ghamizi2022adversarial]. It introduces a task attack rate to optimally weigh each task’s gradient following the loss balancing strategy from GradNorm [chen2018gradnorm], a multi-task optimization method.
4.2 Attack on Clean Multi-Task Models
| Model Index | #Params (M) | PGD | APGD | Auto-SAGA | |||||||||||
| Segm | Norm | Dept | Total | SignTotal | DGBA | Segm | Norm | Dept | Total | SignTotal | DGBA | Baseline | DGBA | ||
@@ 4.2 Attack on Clean Multi-Task Models
| Model Index | #Params (M) | PGD | APGD | Auto-SAGA | |||||||||||||
| Segm | Norm | Dept | Keyp | Edge | Total | DGBA | Segm | Norm | Dept | Keyp | Edge | Total | DGBA | Baseline | DGBA | ||
@@
This subsection compares DGBA with baselines on their effectiveness in attacking clean multi-task models. Tables 1 and 2 compare the attack performance of the baselines and DGBA at the model level on NYUv2 and Tiny-Taskonomy respectively at . Overall, DGBA integrated with PGD and APGD have the highest ARP in 7 out of the 8 models on NYUv2, and in 6 out of the 8 models on Tiny-Taskonomy. Full tables for all models and more results with are included in E.2.
@@
DGBA outperforms baselines Total, SignTotal, and Single-x in attack performance because it alleviates the baselines’ limitations discussed in Attack Framework. As illustrated in Figure 1, Single-x achieves limited attack effectiveness on tasks that are not the attack target x (also discussed in Section 5); Total shares a similar pattern of attack effectiveness across tasks as Single-Segm due to the issue of gradient dominance, making it less effective in attacking all tasks simultaneously. In contrast, DGBA dynamically balances the attack directions for all tasks, making it more threatening for systems that require high robustness. This is also the reason why DGBA outperforms the multi-model attack approach AutoSAGE — DGBA balances the gradients in AutoSAGE.
@@
Figures 2 and 3 further report the influence of the perturbation bounds on the attack performance with different backbone models, Deeplab-ResNet34 and MobileNet respectively. We vary the maximum perturbation bound from 1 to 16 and report the average of the ARP over all multi-task models with diverse architectures. We denote the average ARP as the overall ARP. DGBA is the best-performing attack in almost all and dataset settings, and when integrated with any of the three existing single-task (i.e., FGSM, PGD, APGD) or the multi-model attack (i.e., Auto-SAGA) approaches. There are a few cases, where for , DGBA and Total converge or perform almost identically. This is because at higher values, the magnitude of the noise becomes larger (and thus more visible) and all attacks become more effective. We provide visual adversarial samples in E.4.
Figure 4 presents the attack performance comparisons between DGBA and WGD, built on top of PGD and APGD attack methods. The left panels report the average ARP over multi-task models trained with Deeplab-ResNet34 on the NYUv2 dataset, while the right figures illustrate the attack effectiveness on each individual task, similar to Figure 1. Overall, DGBA significantly outperforms WGD spanning all settings and achieves more balanced attack effectiveness across all tasks compared to WGD. When generating perturbations, WGD assigns more importance to tasks experiencing larger perturbations during the attack, which could amplify the task dominance phenomenon and thus lead to worse attack performance compared to DGBA.
4.3 Attack on Adversarially Trained Models
@@
A common single task strategy to defend against adversarial attacks is to leverage adversarial training [madry2018towards, bai2021recent, zhang2020attacks]. This defense involves generating adversarial samples and using them as part of the dataset during training. This typically results in reduced model performance on clean inputs, but increased robustness to adversarial attacks. Although adversarial training has been incorporated with multi-task settings in prior work [mao2020multitask, ghamizi2023gat], they typically aim to enhance the robustness of a single primary task by introducing auxiliary tasks. It remains unclear whether adversarial training can simultaneously enhance the robustness of all tasks within a multi-task model, especially when incorporated with our proposed balanced adversarial samples.
| Adv. Train | Single-Segm | Single-Norm | Single-Dept | Total | DGBA |
@@
We adopt the single task Friendly Adversarial Training (FAT) [zhang2020attacks] to MTL. FAT is a PGD-based adversarial training method. In order to apply this defense effectively to MTL, we modify the default PGD to DGBA-PGD to use adversarial examples generated from multiple tasks as opposed to just single task adversarial examples. Our new defense is coined FAT DGBA (see E.4 for the detailed algorithm). We also train multi-task models using FAT with the adapted MTL attacks Single and Total. The accuracy of the models after adversarial training are included in the E.4. After adversarial training, we re-evaluate the multi-task attacks to assess both model robustness and attack effectiveness.
@@
Table 3 reports the ARP of 30 cases, i.e., (without adversarial training + 5 adversarially trained models) 5 attack methods. The multi-task models are trained on NYUv2 with Deeplab-ResNet34 and attacked by PGD-variants. We make two main observations. First, the notable reduction in ARP of attack methods (e.g., from to when attacking with Single-Dept) demonstrates that models enhanced with adversarial training are substantially more robust than the model without such training, especially with FAT Total and DGBA (in blue). Second, from the perspective of attack performance, DGBA is still the most effective attack method (in bold), consistently outperforming the Single-x and Total baselines by up to .
5 Parameter Sharing and Model Robustness
RQ3: How does parameter sharing affect the robustness of multi-task models to adversarial attacks? We raise this question because existing literature in MTL determines the appropriate level of parameter sharing with a focus on optimizing task accuracy, while our study reveals the relationship between parameter sharing and model robustness (measured by ARP) with DGBA. For example, with the level of parameter sharing increases, the ARP increases from 87.58 to 108.57 (see Table 1 DGBA-APGD). This result indicates that a higher degree of parameter sharing is also associated with increased adversarial vulnerability. This increased vulnerability potentially arises from the fundamental trade-off between improved task accuracy from positive task interactions enabled by sharing, and increased susceptibility to adversarial attacks due to the greater transferability of perturbations across shared parameters.
To verify our hypothesis, we first define attack transferability in the MTL context. When using an attack method like Single-x, which attacks only one task in a multi-task model, the rest of the tasks in the multi-task model that are not the target could be affected. We measure attack transferability as the ratio between the performance of the attack on all non-targeted tasks versus the performance of the attack on the targeted task. Mathematically, let x represent a task under attack, and y represent another task. The transferability of attack Single-x is , where represents the number of tasks in the multi-task model and ARP-x represents the average relative performance of task x. If the ratio is close to zero, this means that attacking task x has minimal influence on the performance of task y.
Figure 5 illustrates the relationship between the levels of parameter sharing and attack transferability, showing the results for six multi-task models attacked by APGD Single-x variants. These models represent six levels of parameter sharing, ranging from all-share (AS/5L), where all layers of the backbone model (5L) are shared, to independent models (IND/0L), where no layers are shared. We observe a positive correlation between the degree of parameter sharing and the transferability of attacks in the three Single-x variants. As the level of parameter sharing increases, attack transferability also increases, from to () for Single-Dept, to () for Single-Segm, and to () for Single-Norm. These findings suggest a trade-off in multi-task model design: While sharing more parameters among related tasks enhances task accuracy, it also amplifies attack transferability, thereby reducing model robustness, even when facing single task attacks.
6 Conclusion
@@
This work makes significant contributions in understanding the adversarial robustness of multi-task learning (MTL). We first analyzed adaptions of single task white-box attacks to the MTL domain and experimentally demonstrated their ineffectiveness. We then developed a new attack framework DGBA to effectively attack all tasks in multi-task models. On models trained on the NYUv2 and Tiny-Taskonomy datasets, DGBA achieves the highest overall attack strength and is the strongest attack on 6 out of 8 models for both datasets. From the defense side, we also adversarially trained multi-task models called FAT DGBA. DGBA is the most effective attack, even on these multi-task models. We further analyzed the adversarial transferability of MTL adversarial examples and discovered a new phenomenon: Task sharing can lead to increased adversarial transferability. By highlighting the adversarial vulnerability, our work advocates for future research into developing robust and secure multi-task vision models for safety-critical applications. Another interesting extension of this work is to adapt DGBA to attack large-scale multimodal foundation models—such as GPT-4o or Gemini—which typically operate in black-box or API-limited settings. This presents a non-trivial challenge, as it requires extending DGBA beyond white-box supervision and designing new strategies for gradient-free multi-task attacks.
7 Acknowledgement
@@
Appendix A Notation Table
To improve readability and clarify our use of variables, we summarize the main mathematical symbols used throughout the paper in Table 4. Note that although inputs and perturbations are matrix-valued (e.g., pixel-level images), we follow the common convention in the adversarial learning literature of using italic symbols for both vectors and matrices. This notation is consistent with prior works such as APGD [croce2020reliable] and Auto-SAGA [xu2022securing], and allows for easier comparison with standard formulations in the field.
| Symbol | Type | Role |
| Matrix | Input image | |
| Matrix | Adversarial image | |
| Scalar/Matrix | Ground truth label for task | |
| Scalar func | Loss function for task | |
| Scalar func | Combined multi-task loss | |
| Matrix | Adversarial perturbation | |
| Matrix | Magnitude for perturbation direction | |
| Matrix | Binary mask for perturbation direction | |
| Scalar | Normalized loss increase after applying perturbation | |
| Scalar | Number of tasks in the multi-task model | |
| Scalar | Number of PGD iterations |
Appendix B Related Works
@@
Adversarial Attacks. Adversarial attacks fall into two main categories: white-box and black-box attacks [mahmood2021back]. In white-box attacks, an attacker has access to the target model’s internal information, enabling direct gradient extraction and adversarial example generation [carlini2019evaluating]. On the contrary, in black-box attacks, an attacker has limited model knowledge and uses alternative information sources [chen2017zoo, zhou2020dast] to create adversarial examples. This paper focuses on white-box attacks, as they are generally more effective compared to black-box attacks [croce2020reliable, wang2022iwa].
@@
In recent years, many white-box attack techniques have been developed. Fast Gradient Sign Method (FGSM) [goodfellow2014explaining] generates adversarial examples by introducing non-random noise in the gradient direction of the loss function. Projected Gradient Descent (PGD) [madry2018towards] and Momentum Iterative Method (MIM) [dong2018boosting] improve FGSM by generating adversarial samples in an iterative process. Later, Croce and Hein [croce2020reliable] proposed Auto Projected Gradient Descent (APGD) with adaptive step size and combined it with two complementary attacks [croce2020minimally, andriushchenko2020square], developing the ensemble method APGD that outperforms existing methods on diverse benchmark datasets. In addition to gradient-based strategies, alternative methods have emerged. For instance, the Backward Pass Differentiable Approximation (BPDA) [athalye2018obfuscated] accommodates non-differentiable functions, while the Carlini and Wagner () attack [carlini2017towards] perturbs images with minimal delta to misclassify them.
@@
Multi-Task Learning. In Multi-Task Learning (MTL), researchers develop memory and computation-efficient multi-task models that simultaneously address multiple tasks [ruder2017overview, yao2020survey]. The main challenge lies in determining the parameters to share across tasks to optimize both resource efficiency and task accuracy. This has led to an abundance of multi-task model architectures designed either manually [huang2015cross, jou2016deep, dvornik2017blitznet, kokkinos2017ubernet, ranjan2017hyperface] or automatically [sun2019adashare, zhang2022tree, ahn2019deep, guo2020learning, zhang2022automtl]. This paper focuses on attacking branched multi-task models, which are the most representative in the MTL literature [vandenhende2019branched, guo2020learning, bruggemann2020automated, zhang2022tree]. Besides, branched multi-task models have a wide range of sharing patterns across tasks, facilitating the study of the relationship between the robustness of multi-task models and their sharing patterns.
@@
Robustness of Multi-Task Models. Few studies have examined the robustness of the model in MTL settings. A pioneering study [mao2020multitask] pointed out that the adversarial robustness of deep neural networks increases as the number of tasks increases. Subsequent research [ghamizi2022adversarial, li2023sibling] further emphasizes the importance of selecting suitable tasks for joint learning to create more robust models. While these studies offer intriguing insights, they do not specifically propose adversarial attack methods for general multi-task models. Regarding attacks on multi-task models, MTA [guo2020multi] tries to develop attacks in the MTL setting, however, the generated adversarial samples are task-specific and thus fail to attack all the tasks simultaneously. Some other work [gurulingan2021uninet, sobh2021adversarial] attempted to attack the multi-task model by generating adversarial examples for each image while attacking one task at a time or simply targeting the joint multi-task loss. Recent work [chen2023learning] proposes a task-agnostic adversarial sample generator that can attack multiple independent models trained for different tasks. However, its applicability to the unified multi-task models, where tasks share parameters, is not demonstrated and remains an open question. While SMTA [guo2024stealthy] explicitly targets multi-task models, but with a fundamentally different goal: degrading a single target task, while preserving performance on others, i.e., a stealthy attack. MTADV [wang2024multi] introduces a multi-task adversarial attack framework for facial authentication, but it is designed specifically for cross-modal facial datasets and does not generalize to general vision tasks. Different from existing methods, the proposed DGBA aims at jointly degrading all tasks in a multi-task model, regardless of the model’s sharing structure and target vision tasks.
Appendix C MTL Attack Optimization Approximation